OpenStudy (anonymous):

Simplify the expression: (1/x+3) - (1/x+4) and give your answer in the form of: f(x)/g(x) = ??

OpenStudy (anonymous):

x+4-x-3 =1/ (x+4)(x+3)

OpenStudy (anonymous):

where f(x)=1 and g(x)=(x+4)(x+3)

OpenStudy (anonymous):

Thank you so much, Johnny Breuer!